an electric deipole consists of points charges q1 = +12 nC and q2 = -12nC placed 10 cm apart compute the electric potentials at a point a,b and c


1
Expert's answer
2022-03-02T10:16:10-0500

Answer

Electric field due to any electric dipole is given by

"V=\frac{kPcos\theta}{r}"

On axis V="\frac{Kp}{r}"

Zero electrostatic potential due to an electric dipole at an equatorial point.

Where dipole moment

"P=qd\\=12*10^{-9}*0.1\\=1.2*10^{-9}C.m"